OpenStudy (helpmeplease!!!):

can someone give me 25 easy integer questions using +, -, x and /

OpenStudy (raden):

just 5 from me :) 1) 3 + 5*3 - 8/4 = ... 2) 11 - 2 * 10 : 5+ 28 = ... 3) 3 * 6 + 12 : 4 - 20 = ... 4) 48 : 16 * 2 + 10 - 15 = ... 5) 7 - 14 : 2 + 9 * 3 = ... dont use calculator, yeah :P

OpenStudy (helpmeplease!!!):

thanks

OpenStudy (raden):

you're welcome be happy if u want try to solve :D